# Uncertainty-Guided Optimization on Large Language Model Search Trees

### v2

@@ 5.2 Open-ended generation
We use GPT-2 (Radford et al., 2019) and Llama-2-7b (Touvron et al., 2023) for text generation on articles from the Wikipedia (See et al., 2017), CNN Daily Mail (Hermann et al., 2015), and Reddit TL;DR (Völske et al., 2017) datasets. Since many of the text samples in the Wikipedia dataset end with e.g., references instead of full sentences, we filter for text samples with at least $500$ tokens, resulting in a test set with $332$ token sequences (out of a random subset of originally $1000$ sequences). We use $200$ tokens as input and predict $20$ tokens. We do the same for the CNN Daily Mail dataset, where we end up with $790$ token sequences. We use context length of $300$ and generate $60$ tokens. We also include a summarization task, where the goal is to generate a $40$ token long summary of the input sequence. For this, we use $1000$ random samples from the TL;DR dataset with variable-length contexts. Here, we decode sequences of fixed length instead of stopping at the <EOS> token. We run ULTS with both the Dirichlet prior with $\alpha=10^{-4}$ and the empirical prior.
Figure 4: Decoding experiments with Llama-2-7b and GPT-2.
@@
Figure 4 shows the results. No matter the choice of $k_{\max}$ (and $k$), ULTS yields sequences with the same or a higher log-likelihood while expanding fewer nodes. For example in the summarization task with GPT-2, ULTS with the empirical prior and $k_{\text{max}}=20$ achieves an average log-likelihood of $-18.31$ while expanding only $137.90$ nodes. In contrast, beam search with $k=5$ returns sequences of average log-likelihood $-21.33$ even though expanding more nodes (196 nodes). This is the case for both choices of priors, with the empirical prior encouraging exploration a bit more. The histogram in Fig. 6 (Section C.3) suggests that the distribution $\max_{i}c_{ji}$ is bimodal and not fully captured by the Dirichlet prior. In particular, our choice of Dirichlet prior is slightly too pessimistic. As a result, the search may stop too early and the available budget may not be fully utilized. However, the budget that is used, is used efficiently.
Our recommendation is thus as follows. When efficiency is the main goal, a Dirichlet prior with low concentration is preferable—it performs similarly to beam search with smaller beam widths, while being more efficient. If the search performance is important and an additional tree exploration can be afforded (still more efficient than beam search), then the empirical prior is the best choice—it is also hyperparameter-free.
5.3 Runtime
We analyze the runtime overhead on top of the LLM’s forward pass due to ULTS. Figure 5 shows the wall-clock time per iteration, averaged over all sentences and all iterations, broken down into the time spent on the LLM’s forward pass and the time spent on sampling the optimal values and optimizing the acquisition function. The error bars indicate the $95\%$- confidence intervals. Results are shown for the experiments with $k_{\text{max}}=20$ and the Dirichlet prior—the empirical prior performs similarly since ULTS does not differentiate between them in Alg. 2. We note that the runtime overhead of ULTS is small compared to the time spent to do an LLM forward pass. However, note that our current implementation only expands one node of the search tree in each iteration and thereby only evaluate one token sequence per forward-pass. ULTS can be extended similar like BO can be extended into batch BO. It outside of the present work’s scope but is a promising future work.
Figure 5: Average wall-clock time per iteration.
5.4 Extensions
Method Perplexity $\downarrow$ Diversity $\uparrow$
Beam search 1.16$\pm$0.01 0.33$\pm$0.01
Nucleus sampling 3.69$\pm$0.07 0.82$\pm$0.01
Contrastive search 1.27$\pm$0.01 0.4$\pm$0.01
ULTS 1.16$\pm$0.01 0.33$\pm$0.01
ULTS with $\tilde{u}$ 1.3$\pm$0.01 0.55$\pm$0.01
Table 1: Different utility functions (here, $\tilde{u}$) can be used in ULTS’ acquisition function. E.g., to enforce low perplexity and high diversity.
Let ${\bm{v}}_{i}$ contains the optimal values of the descendants of $x_{i}$ and let $u(x_{c},{\bm{v}}_{i}):=\mathbb{I}[v_{\hat{x}_{c}}=\max_{x_{j}\in\mathrm{children}(x_{i})}v_{\hat{x}_{j}}]$. The acquisition function (1) we employed so far can be seen as a sample-based approximation to the expectation
[…]
From the Bayesian decision-theoretic perspective, $u(x_{c},{\bm{v}}_{i})$ can thus be seen as a utility function, encoding our preferences when making decisions under the posterior belief. Hence, we can replace it with any other utility function just like in standard Bayesian optimization (Garnett, 2023) and then use the posterior samples to derive the acquisition function (Wilson et al., 2018).
For example, we can use a utility function that encodes our preference of picking sequences with minimal repetition, which is desirable to avoid text degeneration (Holtzman et al., 2019). Concretely, we can employ a new utility function that contains a repetition-penalty term: $\tilde{u}(x_{c},{\bm{v}}_{i})=\mathbb{I}[(\log v_{\hat{x}_{c}}+\lambda\,b(\hat{x}_{c}))=\max_{x_{j}\in\mathrm{children}(x_{i})}(\log v_{\hat{x}_{j}}+\lambda\,b(\hat{x}_{j}))]$, where $b(\hat{x}_{c})$ is a log-diversity term (Su et al., 2022). Details in Section C.7. The results of this new utility function is in Table 1: ULTS can achieve good perplexity and diversity at the same time, showing the flexibility of our probabilistic decision-theoretic perspective. More results can be found in Table 3 in Section C.7.
@@ 6 Conclusion
We have discussed our method, ULTS, a probabilistic decision-making algorithm for the decoding process of large-language-model (LLM) search trees. Our method quantifies and leverages computational uncertainty over the maximum value of the optimization process. ULTS exploits the structure of the optimization problem—in particular, the concentration strength of the LLM softmax outputs—by putting a prior over the LLM’s softmax outputs and using the implied posterior samples to guide exploration-exploitation in a non-myopic manner. Our work thus connects two seemingly disjoint fields: Bayesian optimization and natural language processing, opening up opportunities for interesting future work in probabilistic inference for LLMs.
Future work  One can study the effect of using more sophisticated priors, e.g. where iid. is not assumed. One can also consider batched acquisition strategies, similar to batch Bayesian optimization techniques (González et al., 2016; Wu & Frazier, 2016). Moreover, it is also interesting to incorporate the uncertainty over the LLM’s outputs (e.g. in the context of Bayesian LLMs) in order to account for possible miscalibration of the likelihood. Another research direction is to extend the approach to reasoning tasks beyond language tasks, e.g. by including beliefs over external rewards (Zhang et al., 2023), such as ones coming from an RLHF-trained reward model, in addition to the current likelihood rewards. Finally, our work in connecting Bayesian optimization with LLMs paves the way to perform probabilistic reasoning with the tree of thoughts framework (Yao et al., 2023).
Acknowledgements
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting JG. JG thanks Microsoft Research for support through its PhD Scholarship Programme. PH and JG gratefully acknowledge financial support by the DFG Cluster of Excellence “Machine Learning - New Perspectives for Science”, EXC 2064/1, project number 390727645; the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039A); and funds from the Ministry of Science, Research and Arts of the State of Baden-Württemberg. Resources used in this work were provided by the Province of Ontario, the Government of Canada through CIFAR, companies sponsoring the Vector Institute https://vectorinstitute.ai/partners/ and the Natural Sciences and Engineering Council of Canada. AR thanks Apple for support through the Waterloo Apple PhD Fellowship, Natural Sciences and Engineering Council of Canada for its support through the PGS-D program, and the David R. Cheriton Graduate Scholarship.
@@ A.1 Prior assumptions
The symmetry assumption for the Dirichlet prior on the LLM’s softmax outputs likely does not entirely hold—some words occur more often in natural language than others (Kingsley Zipf, 1932). However, ULTS only requires access to the distribution over the maximum of the unexplored part of the search space and not over the arg max. The former is not affected by permutations of the entries in the categorical distributions, which is why we suspect that a symmetric Dirichlet distribution with sufficiently small concentration parameter is a good proxy. Moreover, the empirical prior can also be used to address this limitation, without incurring large overhead.
@@ C.1 On-model experiments
We compare ULTS to beam search on artificially generated search problems from Dirichlet priors. The trees have branching factor $b=8$ and depth $d=5$. Since these trees are so small we optimized the acquisition function in eq. (1) over the the full boundary, i.e. non-recursively. The transition probabilities at each node are sampled from a Dirichlet prior with fixed $\alpha\in\{0.1,0.2,0.5,0.8\}$. The comparison is on-model, i.e., ULTS is run with the ground truth parameter of $\alpha$. We repeat the experiment with different values for the confidence parameter $\varepsilon$ of ULTS from $\{0.05,0.1,0.3\}$. Since the toy problems are so small, the exploration of too many nodes is not an issue and we use $k_{\max}=\infty$. Beam search is run with beam sizes ranging from $1$ to $7$. The results in Fig. 6 show that ULTS dominates across the entire range of hyper-parameters. This suggests that knowledge about concentration strength helps reduce the number of search steps.
@@
C.2 Machine translation
We show some example sequences found with ULTS and beam search. For these examples, we picked sentences where ULTS achieves better BLEU scores. Other examples have mainly the same outputs as beam search while ULTS is more efficient.
@@ C.2 Machine translation
C.3 Different choice of hyperparameter $\alpha$ for the Dirichlet prior
@@
Figure 8: Comparison of two different values $\alpha=0.0001$ and $\alpha=5e-6$ of the Dirichlet prior concentration parameter in ULTS
C.4 Additional runtime results
@@
In Table 2 we show runtime results for beam search with beam size $k=5$ and ULTS with maximal beam size $k_{max}=3$ and $k_{max}=5$. Despite expanding fewer nodes than beam search, ULTS is currently slowerin settings where different nodes expansion in beam search can be batched. However, note that batching is not always possible, e.g. in memeory constraint settings (the memory resources depend on the model size, sequence length, as well as batch size).
C.5 Alternative acquisition function
As discussed in Section 3.4, different selection (and hence backup) strategies can be utilized. Recall that all of our results so far are obtained using the “posterior descendant” strategy defined in (2). Here, we show the corresponding results where the other strategy (“posterior”) is used instead.
First, Fig. 9 shows results under the same setting as in the main text, but both ULTS-Dirichlet and ULTS-Emp use the “posterior” strategy instead. We noticed that this strategy also performs well—it is more efficient than beam search. Moreover, it also achieves better or similar likelihood than beam search in Llama-2-7b.
@@ C.5 Alternative acquisition function
We further compare the “posterior” strategy compared to the “posterior descendant” strategy in Fig. 10. We notice that the “posterior” strategy tends to underexplore compared to the “posterior descendant” strategy. Hence, we use and recommend the “posterior descendant” strategy by default.
@@
C.6 Full results for decoding experiments with LLama-2-7b and GPT-2

### v3

@@ 5.2 Off-model experiments with LLMs
Setup  We set ULTS’ $k_{\text{max}}\in\{2,3,4,5,10,20\}$ and set $\varepsilon$ to a default value of $0.1$. The error bars indicate $\pm 1$ standard error of the mean based on all test sentences. We compare ULTS against beam search, as well as other recent baselines commonly used for LLM decoding: multinomial beam search (Multinomial BS; kool2019stochastic, ), contrastive search (su2022contrastiveframeworkneuraltext, ) nucleus sampling (holtzman2019curious, ), best-of-$k$ sampling (stiennon2020learning, ), speculative decoding (leviathan2023fastinferencetransformersspeculative, ), and DoLA (chuang2024doladecodingcontrastinglayers, ). For beam search, multinomial beam search, and best-of-k baselines, we evaluate beam sizes and numbers of samples $k\in\{1,2,3,4,5,10,20\}$, respectively. We set any other hyperparameters of all baselines as suggested by the original papers or by the transformers library. We use Huggingface’s implementation of the baselines (except for A* which is not part of Huggingface’s repertoire). We use GPT-2 (radford2019language, ) and Llama-2-7b (touvron2023llama2, ) for text generation on articles from the Wikipedia (see-etal-2017-get, ), CNN Daily Mail (hermann2015teaching, ), and Reddit TL;DR (volske2017tldr, ) datasets. [^2] Additional results on machine translation tasks can be found in section C.10. Since many of the text samples in the Wikipedia dataset end with e.g., references instead of full sentences, we filter for text samples with at least $500$ tokens, resulting in a test set with $332$ token sequences (out of a random subset of originally $1000$ sequences). We use $200$ tokens as input and predict $20$ tokens. We do the same for the CNN Daily Mail dataset, where we end up with $790$ token sequences. We use a context length of $300$ and generate $60$ tokens. We also include a summarization task, where the goal is to generate a $40$-token long summary of the input sequence. For this, we use $1000$ random samples from the TL;DR dataset with variable-length contexts. We run ULTS with both the Dirichlet prior with $\alpha=10^{-4}$ and the empirical prior. For more details regarding the choice of prior, see section˜B.2.
Figure 3: Decoding experiments with Llama-2-7b and GPT-2 for text generation on CNN Daily Mail and Wikipedia articles and text summarization for the TL;DR dataset. The methods are evaluated for different computational budgets, i.e. different values of $k$ and $k_{\text{max}}$. ULTS dominates the baselines in low-budget settings. Errorbars indicate two times the standard error of the mean.
@@
Results  Figure 3 shows the results. Baselines that are underperforming are shown instead in fig.˜13 in the Appendix. No matter the choice of $k_{\max}$ (and $k$), ULTS yields sequences with the same or a higher log-likelihood while expanding fewer nodes. This is the case for both choices of priors, with the empirical prior encouraging exploration more. The histogram in fig.˜5 (section˜C.4) suggests that the distribution $\max_{i}c_{ji}$ is bimodal and not fully captured by the Dirichlet prior. In particular, our choice of Dirichlet prior is slightly too pessimistic. As a result, the search may stop too early and the available budget may not be fully utilized. However, the budget that is used, is used efficiently. Our recommendation is thus as follows. When efficiency is the main goal, a Dirichlet prior with low concentration is preferable—it performs similarly to beam search with smaller beam widths, while being more efficient. If the search performance is important and an additional tree exploration can be afforded (still more efficient than beam search), then the empirical prior is the best choice—it is also hyperparameter-free. Despite expanding fewer nodes than beam search, ULTS is currently slower in settings where different nodes expansion in beam search can be batched. However, note that batching is not always possible, e.g. in memory-constrained settings (the memory resources depend on the model size, sequence length, as well as batch size). For more details on the runtime see section˜C.6.
@@ 6 Conclusion
We have discussed ULTS, a probabilistic decision-making algorithm for efficiently finding high-likelihood paths on large, expensive search trees, such as those induced by LLMs. Our method quantifies and leverages computational uncertainty over node values and exploits the structure of log-likelihood rewards by placing a prior over them, using posterior samples to guide non-myopic exploration-exploitation. ULTS combines advantages of heuristic methods like beam search, Monte Carlo tree search, and Bayesian tree search, while crucially supporting backtracking without requiring rollouts, additional models, or complex inference. It efficiently finds high-likelihood sequences while expanding few nodes and maintaining low latency.
Future work  One can explore batched acquisition strategies, similar to batch Bayesian optimization (gonzalez2016batch, ; wu2016parallel, ), or incorporate uncertainty over rewards to account for miscalibration. Finally, our connection between probabilistic inference and LLMs paves the way for probabilistic reasoning in the tree of thoughts framework (yao2024tree, ).
Acknowledgments and Disclosure of Funding
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting JG. JG thanks Microsoft Research for support through its PhD Scholarship Programme. PH and JG gratefully acknowledge financial support by the DFG Cluster of Excellence “Machine Learning - New Perspectives for Science”, EXC 2064/1, project number 390727645; the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039A); and funds from the Ministry of Science, Research and Arts of the State of Baden-Württemberg. PH gratefully acknowledges financial support by the European Research Council through ERC CoG Action 101123955 ANUBIS. Resources used in this work were provided by the Province of Ontario, the Government of Canada through CIFAR, companies sponsoring the Vector Institute https://vectorinstitute.ai/partners/ and the Natural Sciences and Engineering Council of Canada. AR thanks Apple for support through the Waterloo Apple PhD Fellowship, Natural Sciences and Engineering Council of Canada for its support through the PGS-D program, and the David R. Cheriton Graduate Scholarship.
@@ A.1 Prior assumptions
The symmetry assumption for the Dirichlet prior on the LLM’s softmax outputs likely does not entirely hold—some words occur more often in natural language than others [27]. However, ULTS only requires access to the distribution over the maximum of the unexplored part of the search space and not over the arg max. The former is not affected by permutations of the entries in the categorical distributions, which is why we suspect that a symmetric Dirichlet distribution with sufficiently small concentration parameter is a good proxy. Moreover, the empirical prior can also be used to address this limitation, without incurring large overhead.
@@ Appendix B Additional Experimental Details
B.1 Hyperparameters
For beam search, multinomial beam search there are no hyperparameters beyond the beam size $k$. Speculative Search uses greedy search with n-gram based assisted decoding with prompt_lookup_num_tokens = 10. Best-of-k Sampling, we use top-p=0.95. For Nucleus Sampling, we use a temperature parameter of 0.2 and top-p=0.95. For contrastive search, we use penalty parameter $\alpha=0.6$ and top-k=50. For DoLA, we set the number of DoLA layers to "high". We use version 4.38.2. of the transformers package.
B.2 Prior Choice for ULTS
We sampled a set $\mathcal{C}$ of Categorical distributions from the LLM on training sequences for each of the datasets and both of the LLMs. We used $1000$ samples from the training set for each of the datasets and LLMs we considered. They are used for the empirical prior, as well as training data for fitting the concentration parameter $\alpha$ of the Dirichlet prior.
Figure˜4 shows samples for the maximum of Categorical distributions returned by the LLM, as well as the distribution of the maximum of the Categorical distribution from a Dirichlet prior for $\alpha=10^{-1},10^{-4},5\times 10^{-6}$.
For the experiments below, we picked a value for $\alpha$ based on visual overlap of those histogramms. For a comparison between different values of the concentration parameter, please refer to fig.˜7 (section˜C.4). Note that this way of choosing the hyperparameter $\alpha$ is very convenient compared to performing costly cross-validation.
@@ C.1 On-model experiments
We compare ULTS to beam search on artificially generated search problems from Dirichlet priors. The trees have branching factor $B=8$ and depth $D=5$. Since these trees are so small we optimized the acquisition function in eq. (4) over the full boundary, i.e. non-recursively. The transition probabilities at each node are sampled from a Dirichlet prior with fixed $\alpha\in\{0.1,0.2,0.5,0.8\}$. The comparison is on-model, i.e., ULTS is run with the ground truth parameter of $\alpha$. We repeat the experiment with different values for the confidence parameter $\varepsilon$ of ULTS from $\{0.05,0.1,0.3,0.5\}$. Since the toy problems are so small, the exploration of too many nodes is not an issue and we use $k_{\max}=\infty$. Beam search is run with beam sizes ranging from $1$ to $7$. The results in fig.˜5 show that ULTS dominates across the entire range of hyper-parameters. This suggests that knowledge about concentration strength helps reduce the number of search steps.
@@
C.2 ULTS for code completion task
In addition, we test ULTS for code completion using the Deepseek Coder 1.3B LLM [17] and all $164$ sequences from the OpenAI HumanEval dataset [3]. The maximum tree depth is set to $500$ tokens and we stop at the $\langle\text{EOS}\rangle$ token. We use a Dirichlet prior with $\alpha=$5\text{\times}{10}^{-6}$\$. To allow for a task-specific qualitative comparison, we evaluated the generated token sequences w.r.t. the Pass@1 metric (i.e. percentage of test cases passed), see table˜1. ULTS achieves favorable performance in terms of Pass@1 and log-likelihood while expanding fewer nodes and incurring lower latency.
Method Pass@1 (%) $\uparrow$ Log-likelihood $\uparrow$ Node expansions $\downarrow$ Runtime (s) $\downarrow$
Greedy 14.02 -28.000 399.369 $13.557\pm 0.493$
Nucleus sampling 14.63 -28.500 377.671 $10.568\pm 0.429$
Beamsearch-Mult ($k=2$) 27.44 -24.375 496.963 $7.235\pm 0.420$
Beamsearch ($k=2$) 29.27 -22.092 497.390 $8.507\pm 0.520$
A-STAR ($k_{max}=2)$ 34.75 -16.968 267.841 $7.184\pm 0.665$
ULTS ($k_{\text{max}}=2$) 25.00 -16.625 134.835 $4.447\pm 0.421$
ULTS ($k_{\text{max}}=3$) 34.76 -15.500 146.866 $4.936\pm 0.454$
Table 1: Results for the code completion task on HumanEval. ULTS can achieve high performance with minimal latency.
C.3 Machine translation
[…]
Below, we show some example sequences found with ULTS and beam search. For these examples, we picked sentences where ULTS achieves better BLEU scores. Other examples have mainly the same outputs as beam search while ULTS is more efficient.
@@ C.3 Machine translation
C.4 Different choice of hyperparameter $\alpha$ for the Dirichlet prior
@@
Figure 7: Comparison of two different values $\alpha=0.0001$ and $\alpha=5e-6$ of the Dirichlet prior conceAcquisitntration parameter in ULTS
C.5 Number of samples in ULTS
$\bm{N}$ Avg. log-lik node expansions Time in s
1 -9.454 86.64 $3.488\pm 0.187$
10 -9.066 90.75 $3.631\pm 0.208$
100 -9.097 92.28 $3.690\pm 0.211$
1000 -9.074 92.41 $3.718\pm 0.213$
Table 2: Results for different choice of number of samples NN.
Table 2 shows results for an ablation for hyperparameter $N$ on 100 sentences from the machine translation task with maximum beam size $k_{max}=4$. Only for $N=1$ the performance is slightly worse, indicating that a small number like $N=10$ might already be sufficient.
C.6 Runtime
We analyze the runtime overhead on top of the LLM’s forward pass due to ULTS. Figure˜8 shows the wall-clock time per iteration, averaged over all sentences and all iterations, broken down into the time spent on the LLM’s forward pass and the time spent on sampling the optimal values and optimizing the acquisition function. The error bars indicate the $95\%$- confidence intervals. Results are shown for the experiments with $k_{\text{max}}=20$ and the Dirichlet prior—the empirical prior performs similarly since ULTS does not differentiate between them in Alg. 2. We note that the runtime overhead of ULTS is small compared to the time spent to do an LLM forward pass. However, note that our current implementation only expands one node of the search tree in each iteration and thereby only evaluate one token sequence per forward-pass. ULTS can be extended similar like BO can be extended into batch BO. This is outside of the present work’s scope but is a promising direction for future work.
Figure 8: Average wall-clock time per iteration. The overhead induced by ULTS is negligible.
@@ C.6 Runtime
In Table 3 we show total runtime results for beam search with beam size $k=5$ and ULTS with maximal beam size $k_{max}=3$ and $k_{max}=5$. Despite expanding fewer nodes than beam search, ULTS is currently slower in settings where different nodes expansion in beam search can be batched. However, note that batching is not always possible, e.g. in memory-constrained settings (the memory resources depend on the model size, sequence length, as well as batch size).
As a reference regarding the computation of the prior: Building the Dirichlet Prior for a tree of depth 250 an branching size 32256 with 1000 samples on a desktop machine (MacBook M1) with CPU only takes 10:50 min (2.6 secs per level of the tree).
C.7 Alternative acquisition function
As discussed in section˜3.2, different selection (and hence backup) strategies can be utilized. Recall that all of our results so far are obtained using the “posterior descendant” strategy defined in (3). Here, we show the corresponding results where the other strategy (“posterior”) is used instead.
First, fig.˜9 shows results under the same setting as in the main text, but both ULTS-Dirichlet and ULTS-Emp use the “posterior” strategy instead. We noticed that this strategy also performs well—it is more efficient than beam search. Moreover, it also achieves better or similar likelihood than beam search in Llama-2-7b.
@@ C.7 Alternative acquisition function
We further compare the “posterior” strategy compared to the “posterior descendant” strategy in fig.˜12. We notice that the “posterior” strategy tends to under-explore compared to the “posterior descendant” strategy. Hence, we use and recommend the “posterior descendant” strategy by default. In general, the optimal value estimates from the “posterior descendant” strategy are less optimistic due to Jensen inequality. This effect is larger if the optimal values of children lie closer together. We illustrate this with the following two examples: In the first example, one of the children is much better than the other children. In such scenarios there is no big difference between the back-up-ed values. In the second example, where some children are close together in terms over their optimal value estimates, the “posterior descendant” estimate is lower making it more likely for the other siblings to get explored, too. We therefore hypothesize that the overall effect of the “posterior descendant” strategy is that exploration is a bit more encouraged as also observed empirically in Figure 12.
Figure 10: Example 1: Left: “posterior descendant” acquisition function.Right: “posterior” acquisition function.
Figure 11: Example 2: Left: “posterior descendant” acquisition function Right: “posterior” acquisition function.
@@
C.8 Full results for decoding experiments with LLama-2-7b and GPT-2